OpenStudy (anonymous):

A sample of forty rivets from the production line had mean 6.585 and standard deviation 0.148 (both in 1/100 of an inch). On communicating these results to the company headquarters in Europe, a request is made for the summary statistics to be converted into millimetres. Given that one inch is 2.538 centimeters, find the mean and standard deviation of the sample in millimetres.

OpenStudy (anonymous):

I got 167 mm +/- 3.76, but it was wrong.

OpenStudy (kropot72):

1 inch = 2.538cm = 2.538 * 10mm = 25.38mm mean = 6.585 1/100th of an inch = 6.585 * 10^-2 inch mean in mm is found as follows:\[mean=25.38\times 6.585\times 10^{-2}=you\ can\ calculate\]

OpenStudy (anonymous):

why do you multiply the mean with that? I don't get where the 100th of an inch comes in

OpenStudy (anonymous):

isn't it just saying that the mean that is given is to 3 decimal places therefore to 100th of an inch?

OpenStudy (anonymous):

because now the mean is 1.67 which is tiny - it should be in the 100s. which is what i got - 167. but it wasn't right. ><

OpenStudy (anonymous):

OHH wait, i got it. misread the question. you're right, it's 1.67. THANK YOU :) and the standard deviation is 0.0376 :D

OpenStudy (kropot72):

You're welcome :)